Problem 3 A parallel-plate capacitor with plate area 8.60 cm2 and air-gap separation 0.43 mm is connected to a 12.00 V battery, and fully charged. The battery is then disconnected. (a) What is the charge on the capacitor? Submit Answer Tries 0/6 (b) The plates are now pulled to a separation of 0.68 mm. What is the charge on the capacitor now? Submit Answer Tries 0/6 c) What is the potential difference across the plates now? Submit Answer Tries 0/6 d) How much work was required to pull the plates to their new separation? Submit Answer Tries 0/6HELP PLS!!
Answer #1

answer) we know the formula

c=\varepsilonoA/d=8.85*10-12C2/N.m2*860\mum2/430\mum=1.77*10-11F

and for charge

q=CdV=1.77*10-11F*12V=2.124*10-10F

a) so the answer is 2.1*10-10C or 2.12*10-10C or 2.124*10-10C

b) here the answer is same as a) because increase in plate distance does not affect the charge

c) we have q=2.124*10-10C

c1=1.77*10-11*0.43/0.68=1.119*10-11C

so new voltage is

V=q/c1=2.124*10-10F/1.119*10-11c=18.9 V or 19 V

so the answer is 19 V or 18.9 V

d) we have the formula

W=1/2*q*dV=1/2*2.124*10-10*(19-12)=7.434*10-10J

so the answer is 7.4*10-10J or 7.43*10-10J or 7.434*10-10J.